OpenStudy (anonymous):

a 4 kg head of a sledgehammer is moving at 6m/s when it strikes a spike diving it into a log. The duration of the impact is 2ms. Find the time average of the impact force.

OpenStudy (anonymous):

are u solving?

OpenStudy (anonymous):

coz gotta go to sleep its 12:00 in india

OpenStudy (ujjwal):

Force= (change in momentum)/time Here, initial momentum= 4*6=24 kg m/s Final momentum=0 time= 2m sec so, force= (24-0)/(2*10^-3) = 12000 N

OpenStudy (anonymous):

got it ty bro

OpenStudy (ujjwal):

Quick response indeed!